# Selective and efficient quantum state tomography for multi-qubit systems

### v1

@@ Introduction.
The basis transformations $U_{\mathcal{E}}$ and $U_{\mathcal{O}}$ map the computational basis to Greenberger–Horne–Zeilinger (GHZ)-type [70, 71] eigenstates as shown in Eqs. (4)–(5). These transformations are thus the conjugate transpose of the unitary operations $U_{\rm GHZ}$ used to prepare $M$-qubit GHZ states. The potential advantage of measurements in a GHZ-type entangled basis has been demonstrated in the context of reducing the sample complexity for learning Pauli expectation values [72]. These operations consist of a local rotation $R_{x(y)}(\pi/2)$ on one qubit, followed by $M-1$ CNOT gates, as illustrated in Fig. 2, giving a maximum circuit depth that grows linearly with the number of qubits. However, the circuit depth can be reduced to $O(\log M)$ for all-to-all qubit connectivity by parallelizing the CNOT gates [73, 74, 75, 76], as explained in the SM [69].
@@
Figure 3: Experimental implementation. (a) Layout of the five qubits used on the ibm_marrakesh quantum processor. (b) Average reconstruction error $\Delta_{\mathrm{err}}$ for the state $|+y\rangle^{\otimes N}$, as a function of the number of qubits $N$ and the number of CZ operations in the SEEQST circuits. (c) The two quantum circuits for the selective block $S_{18}$ in a 5-qubit state, where $q_{1}$ and $q_{5}$ are off-diagonal. Since they are not directly connected, two additional SWAP operations are required to effectively implement ${\rm CNOT}_{15}$ as required by SEEQST. (d) The two circuits for the selective block $S_{32}$ in a 5-qubit state, where all qubits are off-diagonal.
@@
Despite the presence of additional gate errors in SEEQST circuits, the method offers a significant advantage over standard QST in runtime. Specifically, SEEQST completes a full 5-qubit QST in just $4.5$ minutes, compared to $17.5$ minutes for standard QST. This illustrates a practical trade-off: for many applications, the marginal increase in error from SEEQST may be acceptable, especially since standard QST already contains inherent noise. Thus, SEEQST can greatly reduce experimental overhead while still yielding high-fidelity state estimates.
@@
We have developed a general method, SEEQST, for selective quantum state tomography of specific sets of $2^{N}$ elements within an $N$-qubit density matrix. This method requires only two quantum circuits per set, each employing (depending upon the set) at most $N-1$ CNOT gates and one single-qubit rotation. As we have demonstrated both analytically and in experiments with up to $N=5$ qubits, SEEQST enables efficient determination of specific sets of elements without the need for full tomography, and it can perform full QST using only $12^{N+1}-$ 1 circuits, fewer than the $3^{N}$ circuits in standard QST. To enable easy implementation, we have made our SEEQST code freely accessible [66].
In addition to its selective nature for partial reconstruction and fewer experiments needed for full reconstruction, SEEQST offers another advantage over standard QST methods. Unlike those methods, which require processing the entire data set at once to infer the quantum state, SEEQST enables independent computation of subsets, facilitating parallel and simultaneous estimation for even faster full reconstruction.
@@ Acknowledgments.
0 $2.98\times 10^{-3}$ $8.87\times 10^{-4}$ $2.15\times 10^{-4}$ $7.30\times 10^{-5}$
@@
0 $4.10\times 10^{-3}$ $1.16\times 10^{-3}$ $2.33\times 10^{-4}$ $6.30\times 10^{-5}$
1 $6.92\times 10^{-3}$ $1.41\times 10^{-3}$ $3.05\times 10^{-4}$ $9.20\times 10^{-5}$
2 $1.55\times 10^{-3}$ $4.15\times 10^{-4}$ $1.12\times 10^{-4}$
3 $1.54\times 10^{-3}$ $5.46\times 10^{-4}$ $1.47\times 10^{-4}$
4 $1.54\times 10^{-3}$ $3.97\times 10^{-4}$ $1.24\times 10^{-4}$
5 $4.57\times 10^{-4}$ $1.43\times 10^{-4}$
6 $1.46\times 10^{-4}$
7 $1.58\times 10^{-4}$
8 $1.48\times 10^{-4}$
@@
0 $3.59\times 10^{-3}$ $6.66\times 10^{-4}$ $1.26\times 10^{-4}$ $2.20\times 10^{-5}$
@@
0 $4.15\times 10^{-3}$ $7.23\times 10^{-4}$ $1.70\times 10^{-4}$ $4.60\times 10^{-5}$
1 $4.70\times 10^{-3}$ $7.59\times 10^{-4}$ $1.16\times 10^{-4}$ $3.10\times 10^{-5}$
2 - $2.56\times 10^{-3}$ $1.62\times 10^{-4}$ $3.50\times 10^{-5}$
3 - - $6.32\times 10^{-4}$ $4.30\times 10^{-5}$
4 - $6.36\times 10^{-4}$ $1.43\times 10^{-4}$ $5.40\times 10^{-5}$
5 - - $1.95\times 10^{-4}$ $3.10\times 10^{-5}$
6 - - - $3.80\times 10^{-5}$
7 - - - $3.10\times 10^{-5}$
8 - - - $2.90\times 10^{-5}$

### v2

@@ Introduction.
The basis transformations $U_{\mathcal{E}}$ and $U_{\mathcal{O}}$ map the computational basis to Greenberger–Horne–Zeilinger (GHZ)-type [70, 71] eigenstates as shown in Eqs. (4)–(5). These transformations are thus the conjugate transpose of the unitary operations $U_{\rm GHZ}$ used to prepare $M$-qubit GHZ states. The potential advantage of measurements in a GHZ-type entangled basis has been demonstrated in the context of reducing the sample complexity for learning Pauli expectation values [72]. These operations consist of a local rotation $R_{x(y)}(\pi/2)$ on one qubit, followed by $M-1$ CNOT gates, as illustrated in Fig. 2, giving a maximum circuit depth that grows linearly with the number of qubits. However, the circuit depth can be reduced to $O(\log M)$ for all-to-all qubit connectivity by parallelizing the CNOT gates [73, 74, 75, 76], as explained in the SM [69]. Without entangling gates, SEEQST reduces to a constrained version of standard QST, requiring $3^{N}$ circuits for estimating the full density matrix [69].
@@
Figure 3: Experimental implementation. (a) Layout of the five qubits used on the ibm_marrakesh quantum processor. (b) Average reconstruction error $\Delta_{\mathrm{err}}$ for the state $|+y\rangle^{\otimes N}$, as a function of the number of qubits $N$ and the number of CZ operations in the SEEQST circuits. (c) The two circuits for the selective block $S_{32}$ in a 5-qubit state, where all qubits are off-diagonal. (d) The two quantum circuits for the selective block $S_{18}$ in a 5-qubit state, where $q_{1}$ and $q_{5}$ are off-diagonal. Since they are not directly connected, two additional SWAP operations are required to effectively implement ${\rm CNOT}_{15}$ as required by SEEQST.
@@
Despite the presence of additional gate errors in SEEQST circuits, the method offers a significant advantage over standard QST in runtime. Specifically, SEEQST completes a full 5-qubit QST in just $4.5$ minutes, compared to $17.5$ minutes for standard QST. This illustrates a practical trade-off: for many applications, the marginal increase in error from SEEQST may be acceptable, especially since standard QST already contains inherent noise. Thus, SEEQST can greatly reduce experimental overhead while still yielding high-fidelity state estimates. We further support this in the SM through numerical simulations of SEEQST on full-rank five-qubit states [69].
@@
We have developed a general method, SEEQST, for selective quantum state tomography of specific sets of $2^{N}$ elements within an $N$-qubit density matrix. This method requires only two quantum circuits per set, each employing (depending upon the set) at most $N-1$ CNOT gates and one single-qubit rotation. As we have demonstrated both analytically and in experiments with up to $N=5$ qubits, SEEQST enables efficient determination of specific sets of elements without the need for full tomography, and it can perform full QST using only $12^{N+1}-$ 1 circuits, fewer than the $3^{N}$ circuits in standard QST. Furthermore, instead of processing the entire data set at once to infer the quantum state, as standard QST does, SEEQST enables independent computation of subsets, facilitating parallel and simultaneous estimation for even faster full reconstruction. To enable easy implementation, we have made our SEEQST code freely accessible [66].
A possible extension of SEEQST is integration with threshold-based tomography methods [79]. In such methods, one measures the diagonal elements $\rho_{ii}$ and use these to set a significance threshold $t$. This threshold provides a data-driven criterion to reduce the subsequent measurement process: the two SEEQST experiments for a subset $S_{k}$ are performed only if $\sqrt{\rho_{ii}\rho_{jj}}\geq t$, meaning that at least one of its elements $\rho_{ij}$ is predicted to be non-negligible. This reduces experimental overhead without any a priori assumptions on the state’s structure.
@@ Acknowledgments.
0 $1.19\times 10^{-2}$ $7.09\times 10^{-3}$ $3.43\times 10^{-3}$ $2.33\times 10^{-3}$
@@
0 $1.64\times 10^{-2}$ $9.26\times 10^{-3}$ $3.73\times 10^{-3}$ $2.03\times 10^{-3}$
1 $2.77\times 10^{-2}$ $1.12\times 10^{-2}$ $6.35\times 10^{-3}$ $3.98\times 10^{-3}$
2 $1.24\times 10^{-2}$ $6.86\times 10^{-3}$ $2.96\times 10^{-3}$
3 $1.24\times 10^{-2}$ $8.73\times 10^{-3}$ $3.59\times 10^{-3}$
4 $4.84\times 10^{-3}$ $4.55\times 10^{-3}$
5 $7.37\times 10^{-3}$ $4.70\times 10^{-3}$
6 $5.07\times 10^{-3}$
7 $4.73\times 10^{-3}$
8 $4.68\times 10^{-3}$
@@
0 $1.43\times 10^{-2}$ $8.53\times 10^{-3}$ $2.02\times 10^{-3}$ $7.09\times 10^{-4}$
@@
0 $1.67\times 10^{-2}$ $7.58\times 10^{-3}$ $2.77\times 10^{-3}$ $1.46\times 10^{-3}$
1 $1.88\times 10^{-2}$ $1.12\times 10^{-2}$ $2.83\times 10^{-3}$ $8.95\times 10^{-4}$
2 - $2.04\times 10^{-2}$ $2.85\times 10^{-3}$ $1.32\times 10^{-3}$
3 - - $1.01\times 10^{-2}$ $1.39\times 10^{-3}$
4 - $1.24\times 10^{-2}$ $3.13\times 10^{-3}$ $2.10\times 10^{-3}$
5 - - - $1.62\times 10^{-3}$
6 - - - $1.21\times 10^{-3}$
@@
S3 SEEQST Using Only Single-Qubit Gates
In Fig. 2, we employed two-qubit gates to measure all observables in $\mathcal{E}(\Pi^{S_{k}})$ and $\mathcal{O}(\Pi^{S_{k}})$ in two separate experiments, utilizing the fact that observables within each set commute globally. To implement SEEQST using only single-qubit gates, the relevant observables must commute locally on each qubit. Using Eq. (3), the circuits required to measure a given $\Pi^{S_{k}}$ can be constructed as follows:
$\bigotimes_{l=1}^{N}\begin{cases}I,&\text{if }(i_{l},j_{l})\in\{(0,0),(1,1)\},\\ \{R_{x}\left(\pi/2\right),\ R_{y}\left(\pi/2\right)\},&\text{if }(i_{l},j_{l})\in\{(0,1),(1,0)\}.\end{cases}$ (S11)
If, for a given set, $M$ out of $N$ qubits correspond to odd parity, i.e., $i_{k}\neq j_{k}$, then $2^{M}$ distinct single-qubit rotation circuits are required to fully access the corresponding observables. Consequently, the optimal (minimal) total number of distinct circuits across all $2^{N}$ sets is $3^{N}$ in the general case. Therefore, the SEEQST protocol with only single-qubit gates can be regarded as a constrained variant of standard QST, relying solely on local measurement settings and avoiding entangling operations. This approach has been analyzed in Ref. [63].
We analyzed the estimation error $\Delta_{\mathrm{err}}$ from noiseless simulations of 30 random full-rank density matrices for each $N=1$ to $6$, across varying sample sizes $S$ (number of measurement settings multiplied by number of shots), for both methods. In all cases, the error exhibited the expected $1/\sqrt{S}$ scaling. Additionally, we observed a systematic dependence of the error on the odd-parity index $M$.
For single-qubit-only SEEQST, the error increases with increasing $M$, and the error curves appear nearly equally spaced in log scale across different $M$ values, suggesting exponential dependence on $M$ as shown in Fig. S3.
To capture these features, we adopt the following empirical model for single-qubit-only SEEQST:
$\Delta_{\mathrm{err}}^{\mathrm{SQ}}(N,M)=\frac{2^{A(N)+B(N)\cdot M}}{\sqrt{S}},$ (S12)
where the functions $A(N)$ and $B(N)$ capture the system-size dependence. We found that power-law forms for both were sufficient to fit the observed behavior within the standard deviation of the data (except for the case of high $M$ with small sample sizes):
$A(N)=a_{0}+a_{1}\cdot N^{a_{2}},\quad B(N)=b_{1}\cdot N^{b_{2}},$ (S13)
with fitted coefficients
$a_{0}=-0.9177,\quad a_{1}=-0.24734,\quad a_{2}=1.2529,\quad b_{1}=0.6358,\quad b_{2}=-0.1168.$
In contrast, our SEEQST protocol employing entangling CNOT gates shows no such dependence on $M\geq 1$. Its estimation error is empirically related to the single-qubit model as
$\Delta_{\mathrm{err}}^{\mathrm{CNOT}}(N,M)=\begin{cases}\Delta_{\mathrm{err}}^{\mathrm{SQ}}(N,M=1),&\text{for }M\geq 1,\\ \Delta_{\mathrm{err}}^{\mathrm{SQ}}(N,M=0),&\text{for }M=0.\end{cases}$ (S14)
This highlights the statistical advantage of utilizing entangling gates in SEEQST.
Figure S3: Mean estimation error $\Delta_{\mathrm{err}}$ (with standard deviation shaded) as a function of sample size for the SEEQST protocol on a five-qubit system. The figure demonstrates the statistical advantage of using entangled two-qubit gates compared to implementations only employing single-qubit gates, in terms of the number of copies $S$ of the quantum state $\rho$ required. For SEEQST with two-qubit gates, estimation performance remains consistent across different block parities $M$, except for the diagonal block. In contrast, when restricted to single-qubit gates, the estimation error increases with $M$, indicating degraded scaling. The fitted empirical scaling relation in Eqs. (S12)–(S14) (dashed lines) captures the observed behavior within standard deviation, except for with very small sample sizes for higher $M$.
S4 Numerical simulation
To systematically evaluate the performance of SEEQST in noisy environments, we conducted numerical simulations on full-rank five-qubit states, using two standard noise models: amplitude damping and depolarizing noise. In each case, noise was introduced by applying errors with a given probability or coefficient after every layer of parallel gates in the measurement circuits. All results were obtained using 16,384 shots per experiment, with state reconstruction performed via a gradient-descent algorithm, as described in the main text. The results are summarized in Fig. S4.
We see from the left panels in Fig. S4(a,c) that, in the case of standard SEEQST, the estimation error $\Delta_{\rm err}$ depends more strongly on the parity parameter $M$ under depolarizing noise than in the presence of amplitude damping. From the right panels in Fig. S4(a,c), we further observe that the opposite is the case for $\Delta_{\rm err}$ in single-qubit-only SEEQST; here, the dependence on $M$ is stronger with amplitude damping than with depolarizing noise. Overall, the standard SEEQST protocol is more noise-sensitive than its single-qubit-only variant, highlighting that while entangling circuits offer a statistical advantage, they are also more susceptible to noise. In terms of overall performance, the state fidelity of SEEQST is slightly lower than for standard QST under both amplitude damping and depolarizing noise, as shown in Fig. S4(b,d).
Figure S4: Numerical simulation of SEEQST performance under noise. The effect of (a) amplitude damping and (c) depolarizing noise on the estimation error $\Delta_{\rm err}$ for different SEEQST subsets, categorized by the different parity $M$. Left panels show results for the standard SEEQST protocol, while right panels show the performance for its single-qubit gates only variant. (b, d) Impact of amplitude damping and depolarizing noise on the state fidelity $F$, comparing SEEQST with standard QST. All simulations are for full-rank five-qubit states and 16,384 shots per measurement setting. Solid lines represent the mean over multiple simulation runs, and the shaded regions show the standard deviation.
As described in the main text, we employ two approaches for reconstructing the density matrix: (i) maximum likelihood estimation (MLE) with Cholesky decomposition and gradient-descent optimization over the full parameter space of the density matrix, and (ii) direct estimation, where expectation values $\langle P_{i}\rangle$ of Pauli observables are estimated using measured population frequencies as probabilities for the relevant subset of observables. While direct estimation is significantly more computationally efficient, taking on average $10^{-4}$ seconds per state on a standard quadcore CPU compared to $0.3$ seconds for MLE for five-qubit states, it exhibits critical limitations. Figure S5 illustrates this drawback.
We compute $\Delta_{\rm err}$ across 30 randomly sampled full-rank quantum states for varying numbers of randomly selected subsets (excluding the diagonal subset due to its trivial nature and distinct sample-size scaling). For each number of subsets involved, results are averaged over 10 random selections. The plot shows that while the error for MLE remains nearly constant with increasing subsets due to the enforcement of complete positivity and trace preservation (CPTP), direct estimation matches MLE performance only for a single subset, and deteriorates quickly as more subsets are included. Although this analysis uses simulated data with $16{,}384$ shots, the effect is expected to be stronger in experiments, where additional noise sources exist. Indeed, as we observed in our IBM hardware experiments,direct estimation exhibited deviations from MLE even when applied to a single subset.
Figure S5: $\Delta_{\rm err}$ as a function of the number of non-diagonal Pauli subsets involved in the reconstruction, averaged over 30 randomly generated full-rank 5-qubit states using simulated (16,384 shots per setting). Maximum likelihood estimation with Cholesky decomposition and gradient descent (GD) maintains a stable error across subset counts due to CPTP constraints, whereas direct estimation shows increasing $\Delta_{\rm err}$ with more subsets.